OpenStudy (shadow13):

Arrange the list of numbers in increasing order. 2/5, 4/9, 3/7

OpenStudy (anonymous):

what school r u from?

OpenStudy (shadow13):

k12

OpenStudy (shadow13):

@endlesslight254

OpenStudy (anonymous):

oh cool

OpenStudy (anonymous):

a

OpenStudy (shadow13):

ok

OpenStudy (shadow13):

Find the number halfway between 2/5 and 5/6. (Points : 4)

OpenStudy (shadow13):

hellow need help here

OpenStudy (anonymous):

sorry love >u<

OpenStudy (anonymous):

what r the answer choices?

OpenStudy (shadow13):

its ok

OpenStudy (shadow13):

A. 27/40, B. 2/3, C. 7/10, D. 13/20

OpenStudy (anonymous):

i think its b or c

OpenStudy (shadow13):

ok

OpenStudy (shadow13):

simplify [7sqrt{3}+16sqrt{3}]

OpenStudy (anonymous):

what are the choices??

OpenStudy (shadow13):

\[A. 23\sqrt{3} B. 23\sqrt{6} C \sqrt{26} D \sqrt{29}\]

OpenStudy (shadow13):

help please

OpenStudy (anonymous):

i think its b but idk .3.

OpenStudy (shadow13):

ok